OpenStudy (anonymous):

I have 2 similar right angle triangles, CDE and ADG. The length of CDE's hypotenuse is 1/3 the length of ADG's hypotenuse. The area of CDE is 42. What is the area of ADG?

OpenStudy (anonymous):

lengths are in ratio 1:3 areas are in ratioo 1 : 3^2 = 1: 9 area of adg = 9 * 42 = 378

OpenStudy (anonymous):

Thank you! Can you explain why the areas are in ratio 1 : 3^2 ?

OpenStudy (anonymous):

im afraid i cant explain - ill have to check that out

OpenStudy (anonymous):

How did you know to go to that ratio?

OpenStudy (anonymous):

www.mathopenref.com/similartrianglesareas.html - the smaller triangle had hypotenuse of 1/3 of larger so ratio smaller: bigger was 1:3 the corresponding sides must be used ie in this case the two hypotenuse

OpenStudy (phi):

The 2 legs of a right triangle are the base and the height of the triangle. Area of triangle is (1/2) base * height. Since both are 3 times bigger in the larger triangle, you get 3*3=9 times more area.

OpenStudy (anonymous):

the ratio of two areas would be calculated same way for any two triangles - not just right angled in fact its true for any two similar shapes